Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The PJX5 will reduce costs by squeezing more juice from each plum and doing so in a more efficient manner. Mr. Bensen gave Derek the following information. What is the NPV of the PJX5?

a. The PJX5 will cost $1.53 million fully installed and has a 10 year life. It will be depreciated to a book value of $206,022.00 and sold for that amount in year 10.

b. The Engineering Department spent $36,028.00 researching the various juicers.

c. Portions of the plant floor have been redesigned to accommodate the juicer at a cost of $19,398.00.

d. The PJX5 will reduce operating costs by $439,312.00 per year.

e. CSD’s marginal tax rate is 24.00%.

f. CSD is 67.00% equity-financed.

g. CSD’s 20.00-year, semi-annual pay, 6.19% coupon bond sells for $991.00.

h. CSD’s stock currently has a market value of $21.34 and Mr. Bensen believes the market estimates that dividends will grow at 4.09% forever. Next year’s dividend is projected to be $1.49.


Submit
Answer format: Currency: Round to: 2 decimal places.
Answer #1

Given:

Cash outflow for the year 0:

Cost of Machine $1,530,000 + Research cost $36,028 + Redesigning cost $19,398 =$1585426

Yearly Cash inflow in the form of savings in operating cost $439312

Marginal Tax Rate = 24%

Equity financed = 67% ; Balance debt financed = 33%

Calculation of Cost of capital for discounting purpose:

Cost of Equity = Div1/P+g = 1.49/21.34+0.0409 = 0.0698+0.0409 = 0.1107= 11.07%

Cost of Bond = (I + [RV - SV] / n) / (RV + SV) / 2

I= (1000*6.19%*2) =$123.80 ; RV=$1000; SV=$991

= (123.80+(1000-991)/20)/(1000+991)/2 = 124.25/995.5 = 0.1248 AFTER tax = 0.1248(1-0.24) = 0.0949 = 9.49%

WACC= 67%(0.1107)+33%(0.0949) = 0.0742+0.0313 = 0.1055= 10.55%

Calculation on NPV of the PJX5:

NPV = Present value of Cash inflow - Present value of Cash out flow

Calculation of Cash flow after tax = $439312(1-0.24) = $333877.12

Year CFAT

PV Factor

@ 10.55%

Present Value
0 -1585426.00 1 -1585426.00
1-10 333877.12 6 2003940.55
10 (sale value) 206022.00 0.37 75565.95
Total (NPV) 494080.50

NPV of the PJX5 = $494080.50